White dwarf stars are essentially plasmas of free electrons and free protons



(hydrogen atoms that have been stripped of their electrons). Their densities



are typically 5 × 10^9 kg/m3 (i.e., 10^6 times the average density of Earth).



Their further collapse is prevented by the fact that the electrons are highly



degenerate, that is, all the low-lying states are filled and no two identical



electrons can be forced into the same state.



(a) Estimate the temperature of this system. (Assume nonrelativistic electrons and that p2



f /2m = (3/2)kT .)



(b) Now suppose that the gravity is so strong that the electrons are forced



to combine with the protons, forming neutrons (with the release of a



neutrino). If the temperature remains the same, what would be the density



of this degenerate neutron star?




1
Expert's answer
2021-12-24T10:06:03-0500

Density of star="5*10^{9}kg\/m^3"

Mass of star, "M=N(m+2m_p)\equiv2Nm_p"

Where; m=mass of electron, mp= mass of proton,

Electron density, "n=\frac{N}{v}=\frac{\rho}{2m_p}"

="\frac{5*10^9}{2*1.67*10^{-27}}=1.49*10^{36}"

a) "T=\frac{P^2}{3mk}"

"=(\frac{3n}{8π})^{\frac{2}{3}}\frac{h^2}{3*9.1*10^{-31}*1.38*10^{-23}}"

"=(\frac{3*1.49}{8π})^{\frac{2}{3}}\frac{10^{24}*h^2}{3*9.1*10^{-31}*1.38*10^{-23}}"

"=\frac{0.316*10^{24}*10^{54}}{3*9.1*1.38}*(6.62*10^{-24})^2"

"=3.67*10^9K"

b) "T=\frac{P^2}{3m_nk}"

Where mn= mass of neutron

"T=(\frac{3n}{8π})^{\frac{2}{3}}\frac{h^2}{3*1.67*10^{-27}*1.38*10^{-23}}"

"=\frac{0.316*10^{24}*10^{54}}{3*1.67*1.38}*(6.62*10^{-24})^2"

"=2*10^6K"